OpenStudy (anonymous):

Find the fourth roots of 256(cos 280° + i sin 280°).

Parth (parthkohli):

Use the formula (I forgot what's it called, but):\[(\cos(a) + i\sin(a))^{x} = e^{iax}\]

OpenStudy (anonymous):

Where does the 256 come in?

OpenStudy (shubhamsrg):

EULER's formula !

Parth (parthkohli):

But that's an extended Euler's Formula :-)

OpenStudy (anonymous):

e^i(280) ??

OpenStudy (shubhamsrg):

extended ?

Parth (parthkohli):

Yes!

OpenStudy (shubhamsrg):

yep, e^i(280) seems fine What you now have is 256*(e^(i*280)) You have to find 4th root of this.

Parth (parthkohli):

yup.\[\sqrt[4]{ab} =\sqrt[4]{a} \cdot \sqrt[4]{b}\]

OpenStudy (anonymous):

Ahhh! I'm so confused. How do I find the fourth root of 256*(e^(i*280))?

OpenStudy (shubhamsrg):

What's the non-extended formula ?

Parth (parthkohli):

\[e^{i\phi} = \cos(\phi) + i\sin(\phi)\]

OpenStudy (shubhamsrg):

ohh, I never noticed you had put an x there. Sorry about that.

OpenStudy (anonymous):

@shubhamsrg @ParthKohli I really need help. I still do not know how to find the fourth roots of 256(cos 280 + i sin 280)

OpenStudy (anonymous):

I know I have 256*(e^(i*280)), but how do I find the fourth root of that?

Parth (parthkohli):

\[\sqrt[4]{256(\cos280 + i\sin280)} = \sqrt[4]{256}\times \sqrt[4]{\cos(280) + i\sin(280)}\]Remember that \(\rm \sqrt[4]{stuff} = stuff^{1/4}\)?